# Horizons, throats and bounces in hybrid metric-Palatini gravity with a non-zero potential

### v1

@@ I Introduction
The ongoing study of regular black holes Junior:2015fya ; Rodrigues:2015ayd ; Rodrigues:2016fym ; deSousaSilva:2018kkt ; Rodrigues:2019xrc ; Junior:2023ixh ; Junior:2024xmm and black bounces Junior:2022zxo ; Junior:2023qaq ; Fabris:2023opv ; Junior:2024vrv ; Junior:2024cbb within the context of modified theories of gravity is of fundamental importance for advancing our understanding of the spacetime structure and the nature of strong gravitational fields. These studies not only aim to resolve singularities, but also provide crucial insights into the viability of alternative gravity models in describing astrophysical black holes. By investigating deviations from the standard Kerr and Schwarzschild solutions, one may explore how modified gravity can influence event horizon dynamics, causal structures, and the potential observational signatures of non-singular black hole candidates. Beyond the singularity issue, the stability properties of black hole spacetimes within scalar-tensor theories have been a subject of extensive scrutiny Bronnikov:1999wh ; Bronnikov:2024uyb . Stability analyses play a key role in determining whether these black holes are physically viable and whether they can form dynamically from realistic initial conditions. These investigations further elucidate how modifications to the gravitational action impact the intricate interplay between spacetime curvature and matter fields Lobo:2020ffi ; Rois:2024qzm , offering a deeper theoretical framework for understanding deviations from GR. More specifically, hybrid metric-Palatini gravity (HMPG), an extended framework that incorporates both metric and Palatini variational principles Harko:2011nh ; Capozziello:2012ny ; Carloni:2015bua ; Capozziello:2015lza ; Harko:2020ibn ; Lobo:2023ddb ; Capozziello:2012qt ; Capozziello:2013uya ; Capozziello:2013yha , has emerged as a compelling approach for investigating non-singular black hole solutions. This formulation naturally modifies the Einstein field equations by introducing new dynamical terms that arise from the simultaneous variation of the metric and connection, leading to enriched phenomenology in the strong-field regime.
@@ II.1 Action: Jordan frame
It is possible to rewrite the action (1) in an equivalent scalar-tensor representation given as (we refer the reader to Refs. Harko:2011nh ; Capozziello:2012ny ; Danila:2018xya for specific details):
@@ III.3 Horizons, throats and bounces
However, it is also possible for a throat to not be associated with a wormhole geometry. This occurs, for example, in cases where an anti-throat is present Rodrigues:2022mdm , as such configurations do not support a second spatial infinity. Another notable example occurs whenever the minimum of $R(x)$, independently of the sign of $g_{00}$ of the region where it is located, is inside one or more horizons, whichever they are. This geometry is known as a “black bounce spacetime”, or simply “black bounce” Simpson:2018tsi . Essentially, it represents a type of regular black hole where the central singularity is removed and replaced by a regular minimum of $R(x)$. Although the term suggests the existence of a bounce, as defined before, it is used to refer to any spacetime featuring a minimum of $R(x)$ in such a configuration. When this minimum is followed by an expanding universe, the geometry is also known as “black universe”. More generally, this term is used to describe spacetimes where the region beyond the horizon leads to an expanding cosmological solution rather than a singularity.
@@ IV.1 Metric
In this first example, in order to solve the Einstein-frame field equations (10) and (11), we start by assuming, for the spherical radius the following function (as is done in Bronnikov:2020vgg ):
@@
Here, $M$ represents the mass of the configuration, which corresponds to the Schwarzschild mass of our Einstein-frame solution. Based on the preceding discussion, we have the following:
@@
where $\bar{\phi}_{0}\coloneqq\bar{\phi}(\infty)$ is an integration constant that can take any arbitrary value. From these equations, it is evident that all the functions are well-defined within the domain $x\in]-\infty,-1[\;\cup\;]0,+\infty[$, and they diverge at $x=-1$ and $x=0$. These divergences indicate potential singular points. Additionally, the scalar field potential is negative throughout this domain, tending to $\infty$- ∞ as $x$ approaches either $x=-1$ or $x=0$, irrespective of the values of the constants. Conversely, the behaviour of $A(x)$, particularly whether it possesses zeros and the number of such zeros, will depend on the relationship between $M$ and $q$, as will be demonstrated in the following analysis.
@@
Now, we finally have the conditions necessary to analyse the Jordan-frame solution. First of all, by analysing Eq. (29), we see that, as was discussed before, the largest domain of validity of the coordinate $x$ for which the line element takes real values is $x\in\,]-\infty,-1[\;\cup\;]0,+\infty[$. This way, as the metric has to be defined in a continuous interval, there are two distinct intervals of validity: $x\in\,]-\infty,-1[$ or $x\in\,]0,+\infty[$. Note that these are two possible ranges of $x$ that are not simultaneously valid. In fact, by analysing the metric, we find that both show similar behaviour, due to the freedom of choice of the constants, thus, we only analyse the positive one.
@@ IV.3 Characterization of the horizons
Considering these results, we conclude that this solution is analogous to the Reissner-Nordström solution. Consequently, the Penrose diagrams of the spacetime geometries analysed, shown in Figs. 2 and 3, are also analogous to those of that solution. Figure 2 refers to the case $M<M_{c}$, the plots from left to right in Fig. 3 correspond to the diagrams of the cases $M=M_{c}$ and $M>M_{c}$, respectively. More specifically:
@@ V.1 Metric
In this second example, in order to solve the Einstein-frame field equations (10) and (11), we assume a different expression for the spherical radius, given by (as is done in Bronnikov:2020vgg ):
@@
By considering these assumptions and requirements, and following the same procedure that allowed us to obtain Eqs. (25)–(27), we get the following general solution for the current case:
@@
$U(\bar{\phi}(x))$ $=$ $\frac{1}{2\left(x^{2}+1\right)^{2}}\left\{4\left(x^{2}+1\right)\arctan(x)\left[M\left(9x^{2}+3\right)-\left[q^{2}\left(3\pi x^{2}-6x+\pi\right)\right]+q^{2}\left(3x^{2}+1\right)\arctan(x)\right]\right.$ (37)
@@
where $\bar{\phi}_{0}\coloneqq\bar{\phi}(0)$ is an integration constant, and $M$ represents the mass parameter of the Schwarzschild metric. Notably, $M$ also corresponds to the Schwarzschild mass of the present solution in the Einstein frame. Upon analysing these functions, we observe that they are well-defined for all real numbers, remaining regular throughout the domain $x\in]-\infty,+\infty[$, as no divergences occur. Furthermore, the characteristics of the scalar field potential, particularly its sign and the presence of extrema, as well as the behaviour of $A(x)$, including the presence and number of zeros, will depend on the relationship between $M$ and $q$. These dependencies will become evident in the subsequent analysis.
@@
Following the same type of analysis carried out in the previous example, we now start analysing this solution. Firstly, by considering Eq. (40), we find that the largest interval of validity of the coordinate $x$, for which the metric takes real values, is $x\in\,]-\infty,+\infty[$. In general, at both endpoints the metric is regular and $R(x)\to\infty$ (the spherical radius of Eq. (40), which involves the conformal factor), being these the only points that may correspond to a spatial infinity. There are, however, certain scenarios in which the conformal factor prevents $R(x)$ from approaching infinity, thereby also preventing a spatial infinity from arising at both limits simultaneously, as we will see.
If we treat $x\to\infty$ to be the (first) spatial infinity, when it can be considered as such, then $x$ is restricted to the range $-\infty\leq x_{\text{min}}<x<\infty$, where $x_{\text{min}}$ represents the same as before (even though the Einstein frame metric is everywhere regular and its radius never null, this may be a singular or regular central point, due to the conformal factor, as we will see). On the other hand, if we treat $x\to-\infty$ to be the (first) spatial infinity, $x$ is then restricted to the range $-\infty<x<x_{\text{max}}\leq\infty$, where $x_{\text{max}}$ is analogous to $x_{\text{min}}$.
@@ VI Conclusion
These findings emphasize that, with suitable parameter choices, the inclusion of a non-zero scalar potential significantly extends the variety of geometric and causal structures beyond those found in earlier HMPG studies that assumed a vanishing potential. This enrichment manifests in the emergence of novel horizon configurations, throat structures, and regular black hole alternatives, demonstrating the profound impact of the scalar sector on the theory’s solution space. By broadening the range of possible solutions, this work not only enhances the astrophysical relevance of HMPG but also uncovers its potential to describe entirely new families of black hole and wormhole geometries. These findings could have important implications for gravitational physics, guiding future research into the stability properties of these solutions, their observational signatures, such as gravitational lensing and gravitational wave emission, and their connections to quantum gravity and beyond-standard-model extensions of General Relativity.

### v2

@@ I Introduction
The ongoing study of regular black holes Junior:2015fya ; Rodrigues:2015ayd ; Rodrigues:2016fym ; deSousaSilva:2018kkt ; Rodrigues:2019xrc ; Junior:2023ixh ; Junior:2024xmm and black bounces Junior:2022zxo ; Junior:2023qaq ; Fabris:2023opv ; Junior:2024vrv ; Junior:2024cbb within the context of modified theories of gravity is of fundamental importance for advancing our understanding of the spacetime structure and the nature of strong gravitational fields. These studies not only aim to resolve singularities, but also provide crucial insights into the viability of alternative gravity models in describing astrophysical black holes. By investigating deviations from the standard Kerr and Schwarzschild solutions, one may explore how modified gravity can influence event horizon dynamics, causal structures, and the potential observational signatures of non-singular black hole candidates. Beyond the singularity issue, the stability properties of black hole spacetimes within scalar-tensor theories have been a subject of extensive scrutiny Bronnikov:1999wh ; Bronnikov:2024uyb . Stability analyses play a key role in determining whether these black holes are physically viable and whether they can form dynamically from realistic initial conditions. These investigations further elucidate how modifications to the gravitational action impact the intricate interplay between spacetime curvature and matter fields Lobo:2020ffi ; Rois:2024qzm , offering a deeper theoretical framework for understanding deviations from GR. More specifically, hybrid metric-Palatini gravity (HMPG), an extended framework that incorporates both metric and Palatini variational principles Harko:2011nh ; Capozziello:2012ny ; Carloni:2015bua ; Capozziello:2015lza ; Harko:2020ibn ; Lobo:2023ddb ; Capozziello:2012qt ; Capozziello:2013uya ; Capozziello:2013yha ; Rosa:2018jwp ; Rosa:2021yym ; Rosa:2020uoi , has emerged as a compelling approach for investigating non-singular black hole solutions. This formulation naturally modifies the Einstein field equations by introducing new dynamical terms that arise from the simultaneous variation of the metric and connection, leading to enriched phenomenology in the strong-field regime.
@@ II.1 Action: Jordan frame
Now, by introducing a new auxiliary field $E$, the hybrid metric-Palatini action (1) can be reformulated in the equivalent form of a scalar-tensor theory, given by
$S=\int\mathrm{d}^{4}x\sqrt{-g}[R+f(E)+f^{\prime}(E)(\mathcal{R}-E)]+S_{m},$ (3)
where for $E=\mathcal{R}$, the action (3) reduces to action (1). By introducing the following definitions:
$\phi\equiv f^{\prime}(E)\,,\qquad V(\phi)=Ef^{\prime}(E)-f(E)\,\,,$ (4)
it is possible to reformulate Eq. (1) into a scalar-tensor theory with the following action Harko:2011nh :
$S=\int d^{4}x\sqrt{-g}[R+\phi\mathcal{R}-V(\phi)]+S_{m}\,\,.$ (5)
Then, by varying the action (5) with respect to the metric $g_{\mu\nu}$, the scalar field $\phi$ and the independent connection $\hat{\Gamma}^{\alpha}_{\mu\nu}$, respectively, we obtain the field equations
$R_{\mu\nu}+\phi\mathcal{R}_{\mu\nu}-\frac{1}{2}(R+\phi\mathcal{R}-V)g_{\mu\nu}$ $=\kappa^{2}T_{\mu\nu}\;,$ (6)
$\mathcal{R}-V^{\prime}(\phi)$ $=0\;,$ (7)
^$\hat{\nabla}_{\alpha}\big{(}\sqrt{-g}\phi g^{\mu\nu}\big{)}$ $=0\;,$ (8)
where $V^{\prime}(\phi)\equiv dV(\phi)/d\phi$. Moreover, it is possible to relate $\mathcal{R}_{\mu\nu}$ and $R_{\mu\nu}$ using
$\mathcal{R}_{\mu\nu}=R_{\mu\nu}+\frac{3}{2\phi^{2}}\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{\phi}\left(\nabla_{\mu}\nabla_{\nu}\phi+\frac{1}{2}g_{\mu\nu}\Box\phi\right)\,,$ (9)
and rewrite the action (1) in an equivalent scalar-tensor representation given as (we refer the reader to Refs. Harko:2011nh ; Capozziello:2012ny ; Danila:2018xya for more details):
@@ III.3 Horizons, throats and bounces
However, it is also possible for a throat to not be associated with a wormhole geometry. This occurs, for example, in cases where an anti-throat is present Rodrigues:2022mdm , as such configurations do not support a second spatial infinity. Another notable example occurs whenever the minimum of $R(x)$, independently of the sign of $g_{00}$ of the region where it is located, is inside one or more horizons, whichever they are. This geometry is known as a “black bounce spacetime”, or simply “black bounce” Simpson:2018tsi . Essentially, it represents a type of regular black hole in which the central singularity is removed and replaced by a regular minimum of $R(x)$, beyond which the interior evolves into an expanding cosmological solution. Although the term suggests the existence of a bounce, as defined before, it is used to refer to any spacetime featuring a minimum of $R(x)$ in such a configuration. When one of the infinities is flat or AdS and the other is dS, this geometry is also known as “black universe” Bolokhov:2012kn ; Bronnikov:2006fu .
@@ IV.1 Metric
In this first example, in order to solve the Einstein-frame field equations (17) and (18), we start by assuming, for the spherical radius the following function (as is done in Refs. Bronnikov:2020vgg ; Zloshchastiev:2004ny ):
@@
Here, $M$ represents the mass of the configuration, which corresponds to the Schwarzschild mass of our Einstein-frame solution. Based on the preceding discussion, we have the following solution in the Einstein frame:
@@
where $\bar{\phi}_{0}\coloneqq\bar{\phi}(\infty)$ is an integration constant that can take any arbitrary value. It is worth noting that the solution for the particular case $q=0$ has been obtained and discussed in Refs. Bronnikov:2020vgg ; Zloshchastiev:2004ny .
From these equations, it is evident that all the functions are well-defined within the domain $x\in]-\infty,-1[\;\cup\;]0,+\infty[$, and they diverge at $x=-1$ and $x=0$. These divergences indicate potential singular points. Additionally, the scalar field potential is negative throughout this domain, tending to $\infty$- ∞ as $x$ approaches either $x=-1$ or $x=0$, irrespective of the values of the constants. Conversely, the behaviour of $A(x)$, particularly whether it possesses zeros and the number of such zeros, will depend on the relationship between $M$ and $q$, as will be demonstrated in the following analysis.
@@
Note that the particular case $q=0$ of this solution has also been briefly discussed in Ref. Bronnikov:2020vgg .
Now, we finally have the conditions necessary to analyse the Jordan-frame general solution. First of all, by analysing Eq. (36), we see that, as was discussed before, the largest domain of validity of the coordinate $x$ for which the line element takes real values is $x\in\,]-\infty,-1[\;\cup\;]0,+\infty[$. This way, as the metric has to be defined in a continuous interval, there are two distinct intervals of validity: $x\in\,]-\infty,-1[$ or $x\in\,]0,+\infty[$. Note that these are two possible ranges of $x$ that are not simultaneously valid. In fact, by analysing the metric, we find that both show similar behaviour, due to the freedom of choice of the constants, thus, we only analyse the positive one.
@@ IV.3 Characterization of the horizons
In Appendix A, we present the detailed derivation of the $f(\mathcal{R})$ function from action (1) – which is ultimately obtained in parametric form – corresponding to a solution in the $M>M_{c}$ scenario, in particular with $q=5$, $M=5.50389$ ($M=|M_{c}|+0.5$) and $\psi_{0}=0$.
Considering the results above, we conclude that this solution is analogous to the Reissner-Nordström solution. Consequently, the Penrose diagrams of the spacetime geometries analysed, shown in Figs. 2 and 3, are also analogous to those of that solution. Figure 2 refers to the case $M<M_{c}$, the plots from left to right in Fig. 3 correspond to the diagrams of the cases $M=M_{c}$ and $M>M_{c}$, respectively. More specifically:
@@ V.1 Metric
In this second example, in order to solve the Einstein-frame field equations (17) and (18), we assume a different expression for the spherical radius, given by (as is done in Refs. Bolokhov:2012kn ; Bronnikov:2020vgg ):
@@
By considering these assumptions and requirements, and following the same procedure that allowed us to obtain Eqs. (32)–(34), we get the following general Einstein-frame solution for the current case:
@@
$U(\bar{\phi}(x))$ $=$ $\frac{1}{2\left(x^{2}+1\right)^{2}}\left\{4\left(x^{2}+1\right)\arctan(x)\left[M\left(9x^{2}+3\right)-q^{2}\left(3\pi x^{2}-6x+\pi\right)+q^{2}\left(3x^{2}+1\right)\arctan(x)\right]\right.$ (44)
@@
where $\bar{\phi}_{0}\coloneqq\bar{\phi}(0)$ is an integration constant, and $M$ represents the mass parameter of the Schwarzschild metric. Notably, $M$ also corresponds to the Schwarzschild mass of the present solution in the Einstein frame. It is worth noting that this solution has already been obtained and discussed in Ref. Bolokhov:2012kn , as has the particular case $q=0$ in Ref. Bronnikov:2020vgg .
Upon analysing these functions, we observe that they are well-defined for all real numbers, remaining regular throughout the domain $x\in]-\infty,+\infty[$, as no divergences occur. Furthermore, the characteristics of the scalar field potential, particularly its sign and the presence of extrema, as well as the behaviour of $A(x)$, including the presence and number of zeros, will depend on the relationship between $M$ and $q$. These dependencies will become evident in the subsequent analysis.
@@
Note that the particular case $q=0$ of this solution has also been obtained and discussed in Ref. Bronnikov:2020vgg .
Following the same type of analysis carried out in the previous example, we now start analysing this general Jordan-frame solution. Firstly, by considering Eq. (47), we find that the largest interval of validity of the coordinate $x$, for which the metric takes real values, is $x\in\,]-\infty,+\infty[$. In general, at both endpoints the metric is regular and $R(x)\to\infty$ (the spherical radius of Eq. (47), which involves the conformal factor), being these the only points that may correspond to a spatial infinity. There are, however, certain scenarios in which the conformal factor prevents $R(x)$ from approaching infinity, thereby also preventing a spatial infinity from arising at both limits simultaneously, as we will see.
If we treat $x\to\infty$ to be the (first) spatial infinity, when it can be considered as such, then $x$ is restricted to the range $-\infty\leq x_{\text{min}}<x<\infty$, where $x_{\text{min}}$ represents the same as before (even though the Einstein-frame metric is everywhere regular and its radius never null, this may be a singular or regular central point, due to the conformal factor, as we will see). On the other hand, if we treat $x\to-\infty$ to be the (first) spatial infinity, $x$ is then restricted to the range $-\infty<x<x_{\text{max}}\leq\infty$, where $x_{\text{max}}$ is analogous to $x_{\text{min}}$.
@@ V.3.1 $\psi_{0}\neq\frac{\pi}{2}-\frac{\pi}{2\sqrt{3}}+c_{1}\pi$
In Appendix B, we present the detailed derivation of the $f(\mathcal{R})$ function from action (1) – which is ultimately obtained in parametric form – associated with a particular instance within the family of symmetric black bounce solutions described above, namely with $q=5$, $M=25\pi/3$ ($M=M_{c2}$) and $\psi_{0}=0$.
@@ VI Conclusion
It is worth noting that the Einstein-frame solution we obtain in the phantom sector is the same as the one derived in Ref. Bolokhov:2012kn , also under the assumption (40). In that work a detailed analysis of the solution was carried out, including the classification of all possible geometries – based on the number and nature of the horizons and the asymptotic behaviours – as well as the construction of the corresponding Penrose diagrams. In our work, however, the main focus is not on that solution, but rather on the Jordan-frame version, where, in some cases, the conformal factor introduces significant modifications to the geometry and the global structure of the solutions. In particular, when it vanishes at finite values of the radial coordinate, a singularity emerges, which does not occur in any case in the analysis performed in Ref. Bolokhov:2012kn . On the other hand, when it does not vanish, leading to entirely regular solutions, the results are indeed similar to those of that work, and several geometries are identical. Nevertheless, even in this case, novel geometries emerge due to the conformal factor, which allows the local minimum of $R(x)$ to occur at any real value of $x$, rather than always at $x=0$ as in Ref. Bolokhov:2012kn . In any case, note that the introduction of the conformal factor leads to physically inequivalent solutions, even when the spacetime geometry is identical. In particular, thermodynamic properties such as temperature and entropy differ, implying that the underlying thermodynamic system is fundamentally distinct.
Our findings emphasize that, with suitable parameter choices, the inclusion of a non-zero scalar potential significantly extends the variety of geometric and causal structures beyond those found in earlier HMPG studies that assumed a vanishing potential. This enrichment manifests in the emergence of novel horizon configurations, throat structures, and regular black hole alternatives, demonstrating the profound impact of the scalar sector on the theory’s solution space. By broadening the range of possible solutions, this work not only enhances the astrophysical relevance of HMPG but also uncovers its potential to describe entirely new families of black hole and wormhole geometries. These findings could have important implications for gravitational physics, guiding future research into the stability properties of these solutions, their observational signatures, such as gravitational lensing and gravitational wave emission, and their connections to quantum gravity and beyond-standard-model extensions of General Relativity.